OpenStudy (anonymous):

The functions f(x) and g(x) are described below. f(x) = 32x + 8 g(x) = 32x - 9 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units.

OpenStudy (plainntall):

You might find the answer by evaluating both functions at x=0 and finding the difference

OpenStudy (anonymous):

Ok, so it's the plug 0 in for y and 0 in for x thing right?

OpenStudy (plainntall):

There is no y. What does f(0) =, plug 0 into 32x+8 Then do the same for g(0)=, 32x-9 The subtract g(0) from f(0) to get how many units you shift down. Remember what happens when you subtract a negative integer.

OpenStudy (anonymous):

I ended up w/ f(x) =8, and g(x) = -9.

OpenStudy (plainntall):

You got it, how what is the differnce? What is 8-(-9)?

OpenStudy (anonymous):

If you have two negatives, they turn into a +, so it would be 8 + 9 = 17! But, why am I subtracting 8 from -9? Is it because it's asking me to go down, which means subtract?

OpenStudy (plainntall):

Well it asked how many units you shift it down. At x= 0 it is at 8. For g(x) it is at -9 so you shift it down 8 to get to 0, but you still have to shift it down 9 more unites to get to -9. The total number of unites you have to shift down starting at 8 to get to -9 is 17.

OpenStudy (anonymous):

Oh, okay! I understand, so I need to go down 8, then go down 9. On a graph, it would look like -17, when in reality, it's 17! Ok, I understand thank you! Here's a medal for the great help!

OpenStudy (anonymous):

so the answer is 17 ?

OpenStudy (anonymous):

Can't say, sorry. ;___;